OpenStudy (anonymous):

What is the equation of the given line in standard form? Use integer coefficients.y = 5/7x - 12

OpenStudy (anonymous):

A. -5x + 7y = -12 B. -5x - 7y = -84 C. -5x + 7y = -84 D. 5x + 7y = -84

OpenStudy (anonymous):

i know its either b or c

OpenStudy (anonymous):

Start by multiplying both sides by 7. Write it out and show your work.

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

If you don't show your work, I can't see if you are doing things right or not. We tutors are not allowed just to give or confirm answers. You have to be participating actively. Writing down an answer and asking if it is right is not showing work.

OpenStudy (anonymous):

idk how to show my work for this my mom said c

OpenStudy (anonymous):

Here's a start for you (7)(y) = (7)(5/7x - 12). Can you simplify the left side and expand the right side?

OpenStudy (anonymous):

so it is c

OpenStudy (anonymous):

I encourage you to read the Code of Conduct. In it, you will see why this website is intended for ACTIVE learning because that is the only way you will get the concepts and be able to work problems for yourself eventually. I can give you tips, formulas, suggestions, help organize your thinking and analysis, but you have to do the work. Read the Code. I want to help you, but you have to cooperate.

OpenStudy (anonymous):

ok cool whatever ill just go to yahoo

OpenStudy (anonymous):

i was asking if i was correct your supposed to be able to confirm my answer

OpenStudy (anonymous):

Fine, I'll arrange for a moderator to set you straight on this.

OpenStudy (anonymous):

by tell me if im right or not i mean isnt this what this site is for is to check your answers over and see if you have the right ones